OpenStudy (anonymous):

what value would you add to each side to complete square x^2+18x=-8

OpenStudy (amistre64):

what value do you think we should add?

OpenStudy (anonymous):

I am not sure 9

OpenStudy (anonymous):

x^2+18x = x^2 + 2*9*x +0 here's the formula: a^2 + 2ab + b^2 = (a+b)^2 so, you can see a=x, b=9. you are missing b^2, to complete square you'll need b^2, which is 9^2=81 so add 81 to each side.

OpenStudy (anonymous):

Is That it

OpenStudy (anonymous):

try it, then you tell me the result.

OpenStudy (amistre64):

9 is a good number to use, but it is naked the way it is; square it and were good :)

OpenStudy (anonymous):

I am really confused should I square the complete equation?

OpenStudy (anonymous):

you can't square the whole equation. you should end up with something like this (x+a)^2+b=0 a and be can be positive or negative.

OpenStudy (anonymous):

except for some cases where you can just put them together nicely, of course.

OpenStudy (anonymous):

i had a similar problem and don't understand how to do it and how they came up with the answer it was x^2+16x=-7 the answer was 64

OpenStudy (anonymous):

try to put the left side into this form a^2 + 2ab + b^2. add something to make b^2, remember, you add something, you have to subtract it

OpenStudy (anonymous):

a=1 b=18 c=-8

OpenStudy (anonymous):

I know that I am having problems solving this because I can't find an example to use

OpenStudy (amistre64):

b^2 = (2ab/2a) *b

OpenStudy (anonymous):

ok, I'll do you this one example. x^2+16x=-7 x^2+16x+7=0 x^2+2*x*8+7=0 compare this to a^2 + 2*a*b + b^2=(a+b)^2 you can see clearly that a=x b=8 so if b=8, b^2=64, you need to add 64, but you add 64, you need to subtract 64 then equation becomes x^2 + 2*x*8 + 64 - 64 + 7=0 (a+8)^2 + (-64+7) = 0 (a+8)^2 - 57 = 0 there's the answer

OpenStudy (amistre64):

18/2(1) = 9*9 = 81

OpenStudy (anonymous):

(a+9)^2-73=0 So I would add 81 to both sides

OpenStudy (anonymous):

??? I dont get it.

OpenStudy (anonymous):

x^2+18x=-8 x^2+2*x*9+81-81+8=0 (a+9)^2+(-81+8)=0 (a+9)^2-73=0 I would add 81 to both sides

OpenStudy (anonymous):

exactly. that's right

OpenStudy (anonymous):

Thank you now I have an example to go by

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

5x^2-3=0 no solution is that correct?

OpenStudy (anonymous):

thank you got to go let me know if that is right

OpenStudy (anonymous):

you mean to complete square, no, but there are solutions, add 3 to both sides, then divide both sides by 5, you'll get x^2=3/5=(sqrt(3/5))^2=(-sqrt(3/5))^2 so x=sqrt(3/5) or x=-sqrt(3/5)

OpenStudy (anonymous):

thanks i figured it out square root property

OpenStudy (anonymous):

ah that's good, good luck with ur maths